# SCILLA: SurfaCe Implicit Learning for Large urban Area, a volumetric hybrid solution

### v3

@@ Abstract
Neural implicit surface representation methods have recently shown impressive 3D reconstruction results. However, existing solutions struggle to reconstruct driving scenes due to their large, unbounded, and highly complex nature. Hence, to achieve accurate reconstructions, additional supervision data such as LiDAR, strong geometric priors, and long training times are required. To tackle such limitations, we present SCILLA, a new hybrid implicit surface learning method to reconstruct large driving scenes from 2D images. SCILLA’s hybrid architecture models two separate implicit fields: one representing the volumetric density of the scene, and another one representing the signed distance to the surface. To accurately reconstruct urban outdoor driving scenarios, we introduce a novel volume-rendering strategy that relies on self-supervised probabilistic density estimation to sample points near the surface and transition progressively from volumetric to surface representation. Our solution permits a proper and fast initialization of the signed distance field without relying on any geometric prior on the scene, compared to concurrent methods. By conducting extensive experiments on four outdoor driving datasets, we show that SCILLA can learn an accurate and detailed 3D surface scene representation in various driving scenarios while being two times faster to train compared to previous state-of-the-art solutions.
@@
Figure 1: We introduce SCILLA, a novel hybrid implicit learning method capable to accurately reconstruct large driving scenes from 2D RGB input images solely. Extensive experiments on four popular driving datasets show the superiority of SCILLA’s mesh (left) over previous state-of-the-art methods such as StreetSurf [11] (right). Figure shows both original and textured mesh obtained from both methods.
@@ 1 Introduction
Recently, neural implicit surface representation methods [37, 42] have shown impressive 3D surface reconstruction performance from monocular images, achieving highly detailed and precise mesh representations. Such methods are originally inspired by Neural Radiance Fields (NeRF) [23], and re-framed for multi-view surface reconstruction. NeRF introduced volume rendering for Novel View Synthesis (NVS) by sampling points along camera rays. These points, along with their viewing directions, are fed into a Multi-Layer Perceptron (MLP) to predict their respective colors and densities. The final image is rendered by blending these predicted values using alpha compositing. Conversely, neural surface representation methods aim to approximate the surface using a learnable Signed Distance Function (SDF) [37, 42]. Unlike NeRF, the outputs of the MLP are colors and SDF values instead of colors and volumetric densities, which are later blended using a modified alpha-compositing equation adapted to SDF. An accurately reconstructed surface is represented as a zero-level set of the SDF. Assuming the objects lying at the center of the scene, the SDF is first initialized to form a spherical shape to bound the region of interest. Although current methods achieve impressive reconstruction results, they are mostly applied to simple, bounded and object-centric or landmarks scenes. Applying such methods to unbounded driving scenarios results in sub-optimal reconstruction given their significantly different nature. To overcome such limitations, implicit surface reconstruction of large outdoor driving scenes commonly relies on LiDAR supervision [27, 39] or strong geometric priors [11]. These approaches require longer training time to produce qualitative outdoor scene reconstruction meshes from 2D images.
@@
To address those challenges, we propose SCILLA, a hybrid approach to accurately initialize a signed distance field of an unbounded area and represent complex surfaces from urban outdoor driving scenarios with high accuracy and geometric precision, while requiring half of the training time compared to other state-of-the-art solutions such as StreetSurf [11]. We propose a hybrid architecture that models two separate implicit fields: one representing the volumetric density of the scene and another one representing the signed distance to the surface. SCILLA’s novel volume rendering approach relies on a self-supervised density estimator which permits to sample points near the surface along with a progressive sampling strategy that smoothly initializes and transitions from a volumetric to a surface representation, drastically reducing the model convergence time. Inspired by previous state-of-the-art works [11, 43], we supervise surface normals by relying on normal cues predicted from a pre-trained monocular estimator [8]. We extensively validate our solution on KITTI-360 [20], Pandaset [41], Waymo Open Dataset [32], and nuScenes [3]. By comparing the quality of the extracted meshes from the signed distance field, we show that our method can efficiently reconstruct urban outdoor scenarios showing SoTA accuracy. To summarize, the main contributions of our work are:
  • A novel implicit hybrid learning method to progressively initialize the signed distance field for surface reconstruction on large outdoor driving scenes,
@@
  • A hybrid architecture that models two distinct implicit fields to represent both the volumetric density and the signed distance field,
@@
  • A new volume rendering technique that relies on self-supervised probabilistic density estimation to efficiently sample points near the surface and smoothly transition from a volumetric representation to a surface representation, significantly reducing training times,
  • Various regularization techniques of the signed distance field to enhance our model’s implicit surface representation learning and convergence are presented.
@@ 2 Related work
Traditional surface reconstruction techniques represent a scene either by occupancy grids [7, 26, 34, 35] or relying on point clouds [9, 29, 31, 28]. Occupancy grids-based methods are often limited by the voxel resolution and suffer from high memory consumption. On the other hand, Multi-View Stereo (MVS) techniques [24] generate, from posed images, dense point clouds associated with surface normal that can be utilized in a subsequent surface reconstruction step [16]. These methods heavily depend on the quality of the generated point cloud. Nevertheless, they can accumulate errors due to their complex multi-step pipelines, and are mostly designed for object-centric scenes with high overlaps between the observations. We refer the reader to dedicated surveys for in-depth analysis of such techniques [46, 22, 2].
Neural implicit surface representations.
@@
NeRF-based methods can achieve impressive NVS results but fail to accurately reconstruct scene surfaces given the lack of constraints or priors on the surface when modelling the scene as a density volume. Neural surface reconstruction approaches on the other hand, explicitly represent the surface as an SDF and propose various adaptations to train such methods through differentiable volume rendering [37, 42]. NeuS [37] proposes a novel SDF-based rendering equation that relies on SDF values to compute the alpha blending factor used in volume rendering instead of densities as presented in the original NeRF [23]. Nevertheless, NeuS’s original contribution is designed mainly for object-centric datasets and requires a long training time. Some recent works aim to reduce the training time [38] or reconstruct large-scale highly detailed landmark scenes [19]. However, such approaches are not tailored for large unbounded driving scenarios and still require hours of training. In contrast, our method is built upon NeuS’s formulation and it is designed to reliably represent urban outdoor driving scenes.
Urban outdoor surface reconstruction.
@@
Urban outdoor scene reconstruction is one of the main subjects contributing to resolving many downstream applications, such as inverse rendering and object insertion [39]. Latest SoTA approaches rely on SDF-based surface reconstruction for urban outdoor scenes, to disentangle the scene’s geometry from other scene parameters. Neuralangelo [19], a SoTA method for outdoor landmarks reconstruction, requires many overlapping images and the region of interest of the scene to be bounded to initialize the SDF with a spherical shape. Few methods consider acquisition from a moving camera rig for driving sequences surface reconstruction. Existing solutions are mainly guided by LiDAR such as FEGR [39] and Urban Radianc Fields [27], limiting their large-scale deployment. StreetSurf [11], the most related method to our proposal, presents a solution that disentangles close-range view from distant-range view, alongside a signed distance field initialization technique to match the road surface according to the ego poses of the vehicle. This surface initialization approach is rough, not straightforward, and heavily depends on the ego poses and the assumption that the road and the sidewalk are at the same height in every part of the scene. Additionally, StreetSurf relies on the assumption that self-driving scenes are mostly long and narrow and showcase reconstruction of the scene from only one public dataset (i.e. Waymo Open Dataset [32]). In reality, driving datasets frequently contain scenarios where the scenes are short and curvy, hence, disentangling the close and far range in those situations becomes more challenging. Instead, our solution is more adaptive in comparison to StreetSurf as it does not rely on strong assumptions about the scene’s topology. Furthermore, SCILLA is also more compact and faster to train as the entire scene is represented by a single hybrid implicit field.
Hybrid scene representation.
@@
Some SoTA methods have already investigated the use of hybrid rendering solutions relying on both volumetric and surface composing. Wang et al. [40] introduced a hybrid rendering method based on NeuS [37] that uses a spatially varying kernel, which changes according to the fuzziness of the surface. Based on the kernel, a narrow mesh envelope is extracted around the surface, which then enables ray casting only inside this envelope. This solution helps accelerate both rendering time and quality. Other methods [12, 14, 33] focus on a hybrid representation that both relies on an explicit representation (meshes) and an SDF representation. The hybrid volume-mesh representation enables efficient rendering and less resourceful storage while outputting an implicit geometrical representation. In addition to fast training times, this representation also benefits from easy integration into computer graphics engines. NeuSG [5] combines 3D Gaussian Splatting [17] with NeuS to obtain an accurate Gaussian cloud close to the surface that serves as a direct supervision signal during NeuS training. The aforementioned solutions mainly focus on improving the rendering speed and are better fitted for small scenes with a hybrid representation. In our proposal, we jointly train a hybrid model to benefit from the fast training of a density-based approach along with an accurate geometric representation obtained thanks to the signed distance field learning.
@@ 3 Method
Figure 2: SCILLA overview – Our solution can be divided into two key components: an hybrid scene representation (Sec. 3.2) and a ray-based volumetric rendering that progressively transitions from density to SDF sample alpha composition (Sec. 3.3).
SCILLA models the volumetric density and signed distance field in a hybrid architecture. The details of this architecture are presented in Sec. 3.2, while the novel underlying volume rendering strategy is explained in Sec. 3.3. Our hybrid strategy relies on a self-supervised density distribution estimation to sample points near the surface and progressively transition from a volumetric to an implicit surface representation. To facilitate the convergence of our method, we further introduce regularization techniques for the signed distance field. The overall architecture of our proposed method is illustrated in Fig. 2.
@@ 3.1 Preliminaries
NeRF’s Implicit Volumetric Representation. NeRF [23] takes as input the 3D positions $x_{i}\in\mathbb{R}^{3}$ and the viewing direction $d\in\mathbb{R}^{2}$ of sampled points along each camera ray $r$ and predicts their corresponding densities $\sigma_{i}\in\mathbb{R}$ and colors $c_{i}\in\mathbb{R}^{3}$. This representation of the scene is approximated using a Multi-Layer Perceptron (MLP). Finally, volume rendering is used to alpha-composite the colors of the samples along each ray and render the final pixel color $\hat{C}(r)\in\mathbb{R}^{3}$:
@@
with $\delta_{i}\in\mathbb{R}$ being the distance between the samples along the ray. Note that in SCILLA’s hybrid formulation, we refer to this value as $\alpha^{v}_{i}$ to disambiguate from alpha values derived from the estimated SDF which is introduced in Eq. 3.
SDF representation with NeuS. For an accurate surface-based representation of a 3D scene, NeuS [37] presented a different formulation of the NeRF’s volume rendering equation. They represent the scene as a signed distance function, where the surface is the zero-level set of such function: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. NeuS adoptes a volume rendering formulation introducing a density distribution function which is formulated as the derivative of the sigmoid function $\Phi_{s}(x)=(1+e^{-sx})^{-1}$, that is: $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}$. The standard deviation of $\phi_{s}$ is $1/s$ and, as the network converges to an accurate surface representation, this parameter moves toward zero. Following NeuS’s formulation, the volumetric alpha used in Eq. 1 is replaced by the following equation to approximate the color $\hat{C}(r)$:
@@
3.2 SCILLA’s hybrid architecture
Unlike SoTA methods that singularly represent either a density or an SDF field, our hybrid method SCILLA learns the two different fields simultaneously, as illustrated in Fig. 3. SCILLA expresses the scene using two functions, $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$, encoded respectively by two MLPs. The input 3D position $x$ and viewing direction $d$ are embedded respectively via a positional and directional embedding into $\text{emb}_{x}(x)$ and $\text{emb}_{d}(d)$.
@@ 3.2 SCILLA’s hybrid architecture
$\mathcal{F}_{\Theta}^{c}$ on the other hand takes as input the latent vector $h$ that is outputted by $\mathcal{F}_{\Theta}^{h}$ in addition to the embedded viewing direction $\text{emb}_{d}(d)$ and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF, to predict the color $c=\mathcal{F}_{\Theta}^{c}\left(\text{emb}_{d}(d),h,\overrightarrow{n}\right)$.
Figure 3: SCILLA architecture – Our method uses two MLP functions $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$ to output SDF values and density ($\sigma$ and $f(x)$, respectively) along with color values given an input sample. We inspire from NeRF and design $\mathcal{F}_{\Theta}^{c}$ to output the color given the lattent vector $h$ outputted by $\mathcal{F}_{\Theta}^{h}$, the viewing direction ($d$) and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF.
3.3 SCILLA’s volume rendering
Probabilistic density estimation. Probabilistic density estimation along the ray is employed for a more targeted sampling strategy that samples fewer points and accelerates the overall method’s training time. Guided by the estimated density, $N$ points are sampled along the ray, where higher density indicates the presence of the surface. Using SCILLA’s architecture, alpha values are computed for these samples, and volume rendering is performed to obtain the final color associated with the ray. This resulting color is then compared to the ground truth value to train our model.

### v4

# ViiNeuS: Volumetric Initialization for Implicit Neural Surface reconstruction of urban scenes with limited image overlap

@@ Abstract
Neural implicit surface representation methods have recently shown impressive 3D reconstruction results. However, existing solutions struggle to reconstruct driving scenes due to their large size, highly complex nature and their limited visual observation overlap. Hence, to achieve accurate reconstructions, additional supervision data such as LiDAR, strong geometric priors, and long training times are required. To tackle such limitations, we present  ViiNeuS, a new hybrid implicit surface learning method that efficiently initializes the signed distance field to reconstruct large driving scenes from 2D street view images. ViiNeuS’s hybrid architecture models two separate implicit fields: one representing the volumetric density of the scene, and another one representing the signed distance to the surface. To accurately reconstruct urban outdoor driving scenarios, we introduce a novel volume-rendering strategy that relies on self-supervised probabilistic density estimation to sample points near the surface and transition progressively from volumetric to surface representation. Our solution permits a proper and fast initialization of the signed distance field without relying on any geometric prior on the scene, compared to concurrent methods. By conducting extensive experiments on four outdoor driving datasets, we show that ViiNeuS can learn an accurate and detailed 3D surface representation of various urban scene while being two times faster to train compared to previous state-of-the-art solutions.
@@
Figure 1: We introduce ViiNeuS, a novel SDF initialization method tailored to accurately reconstruct large-scale driving scenes from RGB images with limited overlaps. Extensive experiments on popular driving datasets show the superiority of ViiNeuS’s mesh (left) over previous state-of-the-art methods such as StreetSurf[12] (right). The figure presents both the initial mesh and the textured one as derived from both techniques.
@@ 1 Introduction
Recently, neural implicit surface representation methods [39, 45] have shown impressive 3D surface reconstruction performance from monocular images, achieving highly detailed and precise mesh representations. Such methods are originally inspired by Neural Radiance Fields (NeRF) [24], and re-framed for multi-view surface reconstruction. NeRF introduced volume rendering for Novel View Synthesis (NVS) by sampling points along camera rays. These points, along with their viewing directions, are fed into a Multi-Layer Perceptron (MLP) to predict their respective colors and densities. The final image is rendered by blending these predicted values using alpha compositing. Conversely, neural surface representation methods aim to approximate the surface using a learnable Signed Distance Function (SDF) [39, 45]. Unlike NeRF, the outputs of the MLP are colors and SDF values instead of colors and volumetric densities, which are later blended using a modified alpha-compositing equation adapted to SDF. An accurately reconstructed surface is represented as the zero-level set of the SDF. Assuming objects are centered in the scene, the SDF is initially set to a spherical shape to bound the region of interest. While current methods achieve impressive reconstruction results, they are mainly designed for bounded, high-overlap observations of object-centric or landmark scenes. Applying them to urban scenes with limited image overlap leads to sub-optimal reconstructions due to their significantly different nature. Challenges include partial scene observability (e.g., occlusions from parked cars) and the need to handle both close- and far-range objects. Moreover, the geometry of driving scenes is highly complex, with fine structures and large untextured areas like roads and buildings. To overcome such limitations, implicit surface reconstruction of large outdoor driving scenes commonly relies on LiDAR supervision [29, 41] or strong geometric priors [12]. These approaches require longer training time to produce qualitative outdoor scene reconstruction meshes [20].
@@
To address those challenges, we propose ViiNeuS, a hybrid approach to accurately initialize a signed distance field of a large area and represent complex surfaces from urban outdoor driving scenarios with high accuracy and geometric precision while requiring half of the training time compared to other state-of-the-art solutions [12]. We propose a hybrid architecture that models two separate implicit fields: one representing the volumetric density of the scene and another one representing the signed distance to the surface. ViiNeuS’s novel volume rendering approach relies on a self-supervised density estimator which permits to sample points near the surface along with a progressive sampling strategy that smoothly initializes and transitions from a volumetric to a surface representation, drastically reducing the model convergence time. Inspired by previous state-of-the-art works [12, 46], we supervise surface normals by relying on normal cues predicted from a pre-trained monocular estimator [8]. We extensively validate our solution on KITTI-360 [21], Pandaset [43], Waymo Open Dataset [34], and nuScenes [3]. By comparing the quality of the extracted meshes from the signed distance field, we show that our method can efficiently reconstruct urban outdoor scenarios showing SoTA accuracy.
To summarize, the main contributions of our work are:
  • a novel learning method that progressively initialize a signed distance field from street views with limited images overlap,
@@
  • a hybrid architecture, suited to the complexity of outdoor urban scene, that models two distinct implicit fields to represent both the volumetric density and the signed distance field,
@@
  • a new volume rendering technique that relies on self-supervised probabilistic density estimation to efficiently sample points near the surface, significantly reducing overall training times to obtain a scene mesh,
  • an extensive comparison of SoTA implicit SDF and explicit MVS reconstruction solutions on 4 major driving datasets, showing the adaptability and superiority of ViiNeuS for urban scenes surface reconstruction.
@@ 2 Related work
Multi-view surface reconstruction. Traditional surface reconstruction techniques represent a scene either by occupancy grids [7, 28, 36, 37] or relying on point clouds [9, 31, 33, 30]. Occupancy grids-based methods are often limited by the voxel resolution and suffer from high memory consumption. On the other hand, Multi-View Stereo (MVS) techniques [25] generate, from posed images, dense point clouds associated with surface normal that can be utilized in a subsequent surface reconstruction step [18]. These methods heavily depend on the quality of the generated point cloud. Nevertheless, they can accumulate errors due to their complex multi-step pipelines, and are mostly designed for object-centric scenes with high overlaps between the observations. We refer the reader to dedicated surveys for in-depth analysis of such techniques [50, 23, 2].
@@
Neural implicit surface representations. NeRF-based methods achieve impressive NVS results but fail to accurately reconstruct scene surfaces given the lack of constraints or priors on the surface when modeling the scene as a density volume. Neural surface reconstruction approaches, explicitly represent the surface as an SDF and propose various adaptations to train such methods through differentiable volume rendering [39, 45]. NeuS [39] proposes a novel SDF-based rendering equation that relies on SDF values to compute the alpha blending factor used in volume rendering instead of densities as in NeRF [24]. However, NeuS’s geometric initialization is designed for object-centric datasets and requires long training times. Recent works aim to reduce the training time [40] or reconstruct large-scale highly detailed landmark scenes [20]. Neuralangelo [20], a SoTA method for outdoor landmarks reconstruction, requires many overlapping images assuming the structure to be reconstructed is centered within the scene. Indeed, similar to NeuS, they initialize the SDF with a spherical shape. Such approaches are not tailored for large unbounded driving scenarios and still require hours of training. In contrast, our method is built upon NeuS’s formulation but is designed to reliably represent urban outdoor driving scenes.
@@
Driving scenes surface reconstruction. Few surface reconstruction methods consider acquisition from a moving camera rig for driving sequences surface reconstruction. Existing solution are mainly guided by LiDAR such as FEGR [41] and Urban Radianc Fields [29], limiting their large-scale deployment. StreetSurf [12], the most related method to our proposal, divides the scene into separate close-range and far-range views and uses ego pose-based road surface initialization. This method assumes the road and sidewalk are at the same height across the entire scene, which limits its applicability to long, narrow driving scenarios, as seen in the Waymo Open Dataset [34]. Instead, ViiNeuS does not rely on such topological assumptions, making it more adaptable to a variety of driving scenes, including those that are short and curvy. Additionally, ViiNeuS is more compact and faster to train, as it represents the entire scene with a single hybrid implicit field.
@@
Hybrid scene representation. Some state-of-the-art methods have explored hybrid rendering solutions combining volumetric and surface representations. Wang et al. [42] introduced a hybrid method based on NeuS [39] using a spatially varying kernel that adapts based on surface fuzziness. This kernel allows the extraction of a narrow mesh envelope around the surface, which enables ray casting only inside this envelope, accelerating both rendering time and quality. Other methods [13, 16, 35] combine explicit mesh representations with SDF representations. The hybrid volume-mesh representation reduces storage and resource demands while outputting implicit geometry. Additionally, methods like NeuSG [4] and other 3D Gaussian Splatting-based [19] solutions [11, 15] aim to optimize Gaussians for surface reconstruction, with some approaches directly identifying level sets for surface extraction, such as in Gaussian Opacity Fields [47]. The aforementioned solutions mainly focus on improving the rendering speed and are better fitted for scenes with high image overlap. In our proposal, we jointly train a hybrid model to benefit from the fast training of a density-based approach along with an accurate geometric representation obtained thanks to the signed distance field learning.
@@ 3 Method
Figure 2: ViiNeuS  overview – Our solution can be divided into two key components: an hybrid scene representation (Sec. 3.2) and a ray-based volumetric rendering that progressively transitions from density to SDF sample alpha composition (Sec. 3.3).
ViiNeuS models the volumetric density and signed distance field in a hybrid architecture. The details of this architecture are presented in Sec. 3.2, while the novel underlying volume rendering strategy is explained in Sec. 3.3. Our hybrid strategy relies on a self-supervised density distribution estimation to sample points near the surface and progressively transition from a volumetric to an implicit surface representation from limited overlapping street views. To facilitate the convergence of our method, we further introduce regularization techniques for the signed distance field. The overall architecture of our proposed method is illustrated in Fig. 2.
@@ 3.1 Preliminaries
NeRF’s Implicit Volumetric Representation.
NeRF [24] takes as input the 3D positions $x_{i}\in\mathbb{R}^{3}$ and the viewing direction $d\in\mathbb{R}^{2}$ of sampled points along each camera ray $r$ and predicts their corresponding densities $\sigma_{i}\in\mathbb{R}$ and colors $c_{i}\in\mathbb{R}^{3}$. This representation of the scene is approximated using a Multi-Layer Perceptron (MLP). Finally, volume rendering is used to alpha-composite the colors of the samples along each ray and render the final pixel color $\hat{C}(r)\in\mathbb{R}^{3}$:
@@
with $\delta_{i}\in\mathbb{R}$ being the distance between the samples along the ray. Note that in ViiNeuS’s hybrid formulation, we refer to this value as $\alpha^{v}_{i}$ to disambiguate from alpha values derived from the estimated SDF which is introduced in Eq. 3.
SDF representation with NeuS.
For an accurate surface-based representation of a 3D scene, NeuS [39] presented a different formulation of the NeRF’s volume rendering equation. They represent the scene as a signed distance function, where the surface is the zero-level set of such function: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. NeuS adopts a volume rendering formulation that introduces a density distribution function $\phi_{s}(x)$, which is formulated as the derivative of the sigmoid function $\Phi_{s}(x)=(1+e^{-sx})^{-1}$, that is, $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}$. The standard deviation of $\phi_{s}$ is $1/s$ and, as the network converges to an accurate surface representation, this parameter moves toward zero. Following NeuS’s formulation, the volumetric alpha used in Eq. 1 is replaced by the following equation to approximate the color $\hat{C}(r)$:
@@
SDF field initialization
In [44], the authors show that geometric initialization of the signed distance field (SDF) to a bounded sphere is key for learning an SDF function. However, such initialization assumes full scene observability from the training views and a centered region of interest, which does not apply to driving scenes where the camera mounted on the ego-vehicle remains on the road and follows a linear trajectory. StreetSurf [12] initializes the SDF based on road-surface. Instead, we propose a progressive initialization of the SDF, tailored for the complexities of urban scenes, ensuring effective SDF learning across both close and far-range elements in the entire driving sequence.
3.2 ViiNeuS’s hybrid architecture
Our hybrid method ViiNeuS learns the two different fields simultaneously, as illustrated in Fig 3. ViiNeuS expresses the scene using two functions, $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$, encoded, respectively, by two MLPs. The input 3D position $x$ and the viewing direction $d$ are embedded, respectively, through a positional and directional embedding into $\text{emb}_{x}(x)$ and $\text{emb}_{d}(d)$.
@@ 3.2 ViiNeuS’s hybrid architecture
$\mathcal{F}_{\Theta}^{c}$ on the other hand, takes as input the latent vector $h$ that is outputted by $\mathcal{F}_{\Theta}^{h}$ in addition to the embedded viewing direction $\text{emb}_{d}(d)$ and the normal vector $\overrightarrow{n}$ obtained from the SDF gradient, to predict the color $c=\mathcal{F}_{\Theta}^{c}\left(\text{emb}_{d}(d),h,\overrightarrow{n}\right)$.
Figure 3: ViiNeuS  architecture – Our method uses two MLP functions $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$ to output SDF values and density ($\sigma$ and $f(x)$, respectively) along with color values given an input sample. We inspire from NeRF and design $\mathcal{F}_{\Theta}^{c}$ to output the color given the lattent vector $h$ outputted by $\mathcal{F}_{\Theta}^{h}$, the viewing direction ($d$) and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF.
3.3 ViiNeuS’s volume rendering
Probabilistic density estimation.
Probabilistic density estimation along the ray is employed for a more targeted sampling strategy that samples fewer points and accelerates the overall method’s training. Guided by the estimated density, $N$ points are sampled along the ray, where higher density indicates the presence of the surface. Using ViiNeuS’s architecture, alpha values are computed for these samples, and volume rendering is performed to obtain the final color associated with the ray: